Sunland Company purchased machinery for $861000 on January 1, 2014. Straight-line depreciation has been recorded based on a $50100 salvage value and a 5-year useful life. The machinery was sold on May 1, 2018 at a gain of $14100. How much cash did Sunland receive from the sale of the machinery?

$144120

$122220

$172320

$222420
Answer #1

$172,320

Depreciation = cost - salvage/useful life

= (861000 - 50100)/5

= 810900/5

= $162180

Machine used from Jan 1, 2014 to Jan 1, 2018 i.e, 4 years

  • So, depreciation for these 4 years = 4 * 162180 = $648720
  • Depreciation from in 2018 i.e, used only 4 months

= 162180 * 4/12 = $54060

Therefore total depreciation = 648720 + 54060 = $702780

Book value of asset at the time of sale is

= 861000 - 702780

= $158220

Given that it was sold at a gain of $14100

So, the total cash received at the time of sale

= 158220 + 14100

= $172320